# Agnostic Smoothed Online Learning

### v2

@@ 3.2 Recursive construction of R-Cover for classification
In its simplest form, R-Cover subdivides the horizon $[T]$ into $K$ equal-length epochs and uses a learning with expert advice algorithm on each epoch on a subset of functions from $\mathcal{F}$ that are representative from the data observed in previous epochs. For this simpler version, we can for instance use the classical Hedge algorithm [19] on the projection of $\mathcal{F}$ on the queries observed on previous epochs. While we can show that this algorithm already achieves a sublinear regret (see Section 4.1 and Theorem 8 for a detailed discussion), to achieve a $\approx\sqrt{T}$ regret, we need to use a recursive construction, which we parameterize by a depth parameter $P\geq 0$. Intuitively, the approach mentioned above corresponds to the depth-$1$ algorithm.
@@
 Input: depth $P\geq 0$, start and end times $T_{0}\leq T_{1}$ satisfying $T_{1}-T_{0}\geq 2^{P}$, realizable labeled dataset $S\subset\mathcal{X}\times\{0,1\}$
@@
4if $P=0$ then
5       Fix $f_{S}\in\mathcal{F}$ realizing dataset $S$ and predict $\hat{y}_{t}=f_{S}(x_{t})$ for all $t\in(T_{0},T_{1}]$
6else
7       Fix $f_{S}\in\mathcal{F}$ realizing dataset $S$ and let $T_{1/2}:=\left\lfloor\frac{T_{0}+T_{1}}{2}\right\rfloor$
8      for $\alpha\in\{0,1/2\}$ (epoch $(T_{\alpha},T_{\alpha+1/2}])$ do
9             After iteration $T_{\alpha}$, construct all distinct realizable datasets $S_{1},\ldots,S_{r}\subset\mathcal{X}\times\{0,1\}$ obtained by adding labeled points $(x_{t},y_{t})_{t\in(T_{0},T_{\alpha}]}$ for queries from previous epoch to previous dataset $S$
@@
For the base depth $P=0$, given start and end times $T_{0}<T_{1}$ and a dataset $S$, the algorithm simply selects any arbitrary function $f_{S}\in\mathcal{F}$ that agrees on the query set, that is $f_{S}(x)=y$ for all $(x,y)\in S$, and uses it as prediction at all times in $[T]$.
@@
As an important remark, because $\mathcal{F}$ has VC dimension $d$, we always have $r\leq 2T^{d}+1$ from Sauer-Shelah’s Lemma 1. The additional expert comes from the fact that we also added $f_{S}$ as expert.
@@ 3.3 Learning with expert advice algorithm
Instead of using the standard exponentially weighted algorithm for learning with expert advice, we use a specific variant. We briefly recall the setup for prediction with $K$ experts and fixed horizon $T$ that is relevant for our present discussion. At each iteration $t\in[T]$, the environment chooses losses $\ell_{t,i}$ for each experts $i\in[K]$. The learner then selects an expert $\hat{i}_{t}\in[K]$ potentially randomly without knowledge of the losses at time $t$. Last, all losses at time $t$ are revealed to the learner and they incur the loss $\ell_{t,\hat{i}_{t}}$ from the selected expert. The goal of the learner is to minimize its regret compared to the performance of any fixed expert:
$Reg(T):=\sum_{t=1}^{T}\ell_{t,\hat{i}_{t}}-\min_{i\in[K]}\sum_{t=1}^{T}\ell_{t,i}.$
The classical exponentially weighted forecaster or Hedge algorithm (see e.g. [19]) with parameter $\eta>0$ proceeds as follows. At time $t$, it computes the cumulative regret compared to each expert up to time $t$: $R_{t-1,i}:=\sum_{l=1}^{t-1}\ell_{l,\hat{i}_{l}}-\ell_{l,i}$ for all $i\in[K]$. It then randomly samples $\hat{i}_{t}\sim p_{t}$ where the distribution $p_{t}=(p_{t,i})_{i\in[K]}$ is defined via exponential weights
$p_{t,i}:=\frac{e^{\eta R_{t-1,i}}}{\sum_{j\in[K]}e^{\eta R_{t-1,j}}}.$
We next denote by $\mathcal{F}_{t}=(\ell_{l,i},l\leq t,i\in[K],\hat{i}_{l},l<t)$ the history up to time $t$ included. The exponentially weighted forecaster with learning parameter $\eta$ enjoys the following classical bound (see e.g. [19, Corollary 2.2]):
$PReg(T):=\sum_{t=1}^{T}\mathbb{E}_{\hat{i}_{t}}[\ell_{t,\hat{i}_{t}}\mid\mathcal{F}_{t}]-\min_{i\in[K]}\sum_{t=1}^{T}\ell_{t,i}\leq\frac{\ln K}{\eta}+\frac{T\eta}{2}.$
We will refer to the quantity on the left-hand side as the pseudo-regret $PReg(T)$. Using the standard choice of parameter $\eta=\sqrt{2\ln K/T}$, and assuming that the losses all have values in $[0,1]$, the previous equation directly gives an expected bound on the regret $\mathbb{E}[Reg(T)]\lesssim\sqrt{T\ln K}$. However, for our purposes, we need a refinement of this bound. Using [19, Theorem 2.1], we can derive the following bound
$PReg(T)\leq\frac{\ln K}{\eta}+\frac{\eta}{2}\sum_{t=1}^{T}\sum_{i\in[K]}p_{t,i}r_{t,i}^{2},$ (1)
where $r_{t,i}:=\ell_{t,\hat{i}_{t}}-\ell_{t,i}$ is the instantaneous regret of the forecaster compared to expert $i$ at time $t$. For convenience, let us denote
@@
Eq. 1 yields a tighter bound than the standard regret bound if one selects $\eta\approx\sqrt{\ln K/\Delta_{T}}$ instead of the standard choice $\eta\approx\sqrt{\ln K/T}$. Achieving the corresponding bound without a prior knowledge of $\Delta_{T}$ can be easily performed via the standard doubling trick. Precisely, we use the exponentially weighted forecaster with initial parameter $\eta_{1}\approx\sqrt{2\ln K}$ until $\Delta_{t}\geq 1$, then restart the algorithm with a parameter $\eta_{2}\approx\eta_{1}/2$ until $\Delta_{t}\geq 4$. We continue the process by always restarting the algorithm with a quadrupled threshold for $\Delta$ and a corresponding parameter $\eta>0$ (roughly halved). The precise algorithm is given in Algorithm 2, which is the exponentially weighted forecaster variant that we use for our algorithm R-Cover. This variant enjoys the following pseudo-regret bound, whose proof is given in Appendix B.
@@
3 Let $k=1$, $\Delta_{max,1}=1$, $\eta_{1}=\sqrt{2\ln K/(\Delta_{max,1}+1)}$
4Initialize $R_{0,i}=0$ for all $i\in[K]$, and $\Delta_{1}=0$
@@ 3 Let $k=1$, $\Delta_{max,1}=1$, $\eta_{1}=\sqrt{2\ln K/(\Delta_{max,1}+1)}$
6       Let $p_{t,i}=\frac{e^{\eta_{k}R_{t-1,i}}}{\sum_{j\in[K]}e^{\eta_{k}R_{t-1,j}}}$ and sample $\hat{i}_{t}\sim p_{t}$ independently from history
@@ 11            Reset $R_{t,i}=0$ for all $i\in[K]$, $\Delta_{k+1}=0$, and $k\leftarrow k+1$
Algorithm 2 Adaptive exponentially weighted forecaster A-Exp
@@
Suppose that all losses lie in $[0,1]$. Then, the pseudo-regret of the adaptive exponentially weighted forecaster A-Exp satisfies
$PReg(T)\leq 8\sqrt{\max(\Delta_{T},1)\ln K},\quad T\geq 1.$
@@
$Reg(T)\leq 12\sqrt{\max(\Delta_{T},1)\ln K}+2\ln\frac{1}{\delta}.$
@@ 4 Technical overview
As discussed above, the classification setting will be mostly sufficient to present our main proof ideas. Hence, in this section we mostly focus on this case, that is, we suppose that $\mathcal{F}:\mathcal{X}\to\{0,1\}$ has VC dimension $d$.
@@ 4.1 A simple algorithm for a weaker regret guarantee
To motivate the form of R-Cover, we first consider a significantly simpler algorithm which essentially corresponds to R-Cover with depth $1$. In this simplest form, R-Cover subdivides the horizon $[T]$ into $K$ equal-length epochs and a learning with expert advice algorithm on each epoch on the projection of the function class $\mathcal{F}$ on query points $x_{t}$ from prior epochs. For instance, we can use the classical exponentially weighted forecaster algorithm (e.g. see [19]). This simplified algorithm which we call Cover is summarized in Algorithm 3.
@@
We formally prove this result in Appendix D. In this section, our goal is mostly to give key intuitions about the underlying strategy for the full algorithm R-Cover. To give some insights into why Cover already achieves sublinear regret, note that if the queries prior to some epoch $(T_{k-1},T_{k}]$ are “representative” of the queries during this epoch, then the cover $S_{k}$ constructed at the beginning of the epoch (line 3 of Algorithm 3) is a good representative set of relevant functions. Naturally, this holds if the underlying process $(x_{t})_{t\in[T]}$ is i.i.d.—that is $\sigma=0$. The crux of our analysis is to show that when the adversary is $\sigma$-smooth this still holds in an amortized sense. Note that it is not true that the queries $(x_{t})_{t\leq T_{k-1}}$ observed prior to some epoch $(T_{k-1},T_{k}]$ are always representative of the queries during that epoch. Indeed, a $\sigma$-smooth adversary can for instance decide to have the sequence of distributions $(\mu_{t})_{t\in[T]}$ adaptively switch from one distribution to a completely unrelated one up to $\left\lfloor 1/\sigma\right\rfloor$ times. However, we show that the number of epochs for which prior queries $(x_{t})_{t\leq T_{k-1}}$ are not representative of the queries on the epoch $(T_{k-1},T_{k}]$ is bounded.
To quantify the notion of “representativeness”, we introduce the following quantity, which essentially quantifies the maximum $\ell_{1}$ discrepancy between queries observed until some time $t_{0}<t$ and the query made at time $t$ on the function class $\mathcal{F}$. For any $0\leq t_{0}<t\leq T$, we define
@@
where we recall that $\mathcal{H}_{t-1}$ denotes all history available until the end of iteration $t-1$. Intuitively, if the queries prior to $t_{0}$ were representative of the query at time $t$, then the empirical projection of $\mathcal{F}$ onto the query set $(x_{t})_{t\leq t_{0}}$ should reasonably cover $x_{t}\sim\mu_{t}$ and as a result $\gamma_{t_{0}}(t)$ would be smaller.
One of our main contributions for the analysis of smoothed adversaries is the following result which bounds the number of epochs on which prior history is not representative. How the epochs are constructed is very flexible: we used a fixed schedule for Cover and R-Cover but randomized epochs are also possible, which may be useful for improved regret bounds in the regression case. The proof uses some key results from [15].
@@
Proposition 9 shows that up $\tilde{\mathcal{O}}(1/(q\sigma))$ epochs, we only pay at most a price $w(T,\delta)$ during each epoch $(T_{k-1},T_{k}]$ for the times $t\in(T_{k-1},T_{k}]$ when the cover constructed from queries prior to this epoch was not representative of query $x_{t}$ by some threshold $q$. Here, we largely view $w(T,\delta)$ as a reasonable price to pay on each epoch. Hence, intuitively, we can consider that up to $\tilde{\mathcal{O}}(1/(q\sigma))$ epochs, the cover constructed from queries on prior epochs is always representative from the queries on the epoch up to threshold $q$.
@@
We emphasize that up to the logarithmic factors, Proposition 9 is tight in the following sense. For any choice of the online mechanism to construct epochs and threshold $q$, a $\sigma$-smooth adversary can ensure that for $\mathcal{O}(1/(q\sigma))$ epochs $(T_{k-1},T_{k}]$, queries on prior epochs are not representative up to threshold $q$ from all times in $(T_{k-1},T_{k}]$. We detail below the scenarios for which Proposition 9 is tight. We believe that these essentially captures all possible attack behaviors of a smooth adversary.
@@
Because of the $\sigma$-smoothness constraint, the adversary cannot query the algorithm on completely different regions of the space $\mathcal{X}$ at each epoch. One possible strategy for the adversary, which we discussed as motivation above, is to switch distributions $\left\lfloor 1/\sigma\right\rfloor$ times during the learning process, possibly onto a completely new region of the space. This corresponds to $q=1$ in Proposition 9: at the start of $\left\lfloor 1/\sigma\right\rfloor$ epochs $(T_{k-1},T_{k}]$, the adversary switches query distributions $\mu_{t}$ and selects a distribution with support on a new region for which prior queries are irrelevant. This results in $\gamma_{T_{k-1}}(t)=1$ for all $t\in(T_{k-1},T_{k}]$.
@@
A more refined strategy for the adversary in order to increase its number of affected epochs is to select a parameter $q$ and at the start of a new epoch $(T_{k-1},T_{k}]$, switch the query distribution as follows. They construct a new mixture distribution $\mu_{k}:=q\nu_{k}+(1-q)\mu_{0}$ where with probability $q$ the learner is queried on a new distribution $\nu_{k}$ with say completely new support compared to the history, and with probability $1-q$ the learner is queried on a base measure $\mu_{0}$ that is very similar to previous queries. This results in $\gamma_{T_{k-1}}(t)\geq q$ for all $t\in(T_{k-1},T_{k}]$. On one hand, during the epoch, the adversary could only test the learner on a fraction $q$ of “truly adversarial” queries sampled from $\mu_{k}$. On the other hand, the smoothness constraint is now easier to satisfy and we can check that the adversary can afford to corrupt $\approx 1/(q\sigma)$ epochs in this manner. This precisely corresponds to the bound from Proposition 9 up to logarithmic factors. As it turns out, this mixture strategy is in fact stronger for the adversary and choosing $q\approx 1/\sqrt{T}$ is the strategy that yields the lower bound from Theorem 5.
@@
The statement from Proposition 9 is written specifically for classification, for which analyzing the $\ell_{1}$ diameter as defined in $\gamma_{t_{0}}(t)$ in Eq. 2 is amenable. The proof of Proposition 9 requires controlling the complexity of the class $\{\mathbbm{1}[f\neq g]:f,g\in\mathcal{F}\}$ which has VC dimension bounded by $2d$ if $\mathcal{F}$ has VC dimension $d$. While the VC dimension behaves nicely with this self-difference operation, this is not the case for the fat-shattering dimension which is known to behave somewhat wildly with the addition [1].222[1] notes that the function class $\mathcal{F}$ of increasing functions on $[0,1]$ always has fat-shattering dimension one at any scale, while =$\mathcal{F}-\mathcal{F}=\{f-g:f,g\in\mathcal{F}\}$ has infinite shattering dimension at all scales. To solve this issue for the regression setting, we need to localize this difference class around an oblivious benchmark function $f^{\star}$. The localized analog of $\gamma_{t_{0}}(t)$ that we use in our proofs is defined in Eq. 16. The corresponding generalization of Proposition 9 is Lemma 13. In this general regression setting, the term in $w(T,\delta)$ from Proposition 9 depending on the VC dimension $d$ is replaced by the Wills functional of $\mathcal{F}$ which measures the complexity of the class.
@@
for some constant $C\geq 1$, where we used Jensen’s inequality in the last inequality. Next, for each $k\in[K]$, denote by $f_{k}\in S_{k}$ the function in the cover that had the correct labeling compared to $f^{\star}$, that is:
$f_{k}(x_{t})=f^{\star}(x_{t}),\quad t\in[T_{k-1}],k\in[K].$
Because $f_{k}$ is one of the experts considered during epoch $k$, it suffices to bound the remaining regret term
$\sum_{k\in[K]}\sum_{t=T_{k-1}+1}^{T_{k}}\ell_{t}(f_{k}(x_{t}))-\ell_{t}(f^{\star}(x_{t}))\leq\sum_{k\in[K]}\sum_{t=T_{k-1}+1}^{T_{k}}\mathbbm{1}[f_{k}(x_{t})\neq f^{\star}(x_{t})].$
Taking the expectation of each term for $x_{t}$ conditionally on the history $\mathcal{H}_{t-1}$, we obtain

### v3

@@ 3.1 Recursive construction of R-Cover for classification
In its simplest form, R-Cover divides the horizon $[T]$ in $K$ equal-length epochs and uses a learning with expert advice algorithm on each epoch on a subset of functions from $\mathcal{F}$ that are representative from the data observed in previous epochs. While we can show that this achieves a sublinear regret (see Section 4.1 and Theorem 6 for a detailed discussion), to achieve a $\approx\sqrt{T}$ regret, we need to use a recursive construction, which is parameterized by a depth parameter $P\geq 0$.
@@
 Input: depth $P\geq 0$, times $T_{0}\leq T_{1}$ satisfying $T_{1}-T_{0}\geq 2^{P}$, realizable dataset $S\subset\mathcal{X}\times\{0,1\}$
@@
4Fix $f_{S}\in\mathcal{F}$ realizing dataset $S$, i.e., $f_{S}(x)=y$ for all $(x,y)\in S$. Let $T_{1/2}:=\left\lfloor\frac{T_{0}+T_{1}}{2}\right\rfloor$
5if $P=0$ then
6       for $t\in(T_{0},T_{1}]$ do  Predict $\hat{y}_{t}=f_{S}(x_{t})$;
7
8 else
9       for $t\in(T_{0},T_{1/2}]$ do
@@
13      Construct all distinct realizable datasets $S_{1},\ldots,S_{r}\subset\mathcal{X}\times\{0,1\}$ obtained by adding labeled points $(x_{t},\tilde{y}_{t})_{t\in(T_{0},T_{1/2}]}$ for queries from the previous epoch to $S$, as defined in Eq. 1
14      for $t\in(T_{1/2},T_{1}]$ do
[…]
16       end for
17
@@
For the base depth $P=0$, given start and end times $T_{0}<T_{1}$ and a dataset $S$, the algorithm simply selects any arbitrary function $f_{S}\in\mathcal{F}$ that agrees on the query set, that is $f_{S}(x)=y$ for all $(x,y)\in S$, and uses it as prediction at all times in $[T]$. We recall that $S$ intuitively encodes prior information gathered on the prediction problem, hence, the algorithm for $P=0$ simply corresponds to following this prior information irrespective of the information obtained during the learning process on $(T_{0},T_{1}]$. For larger depths, the algorithm will make use of this information proceeding by epochs.
Suppose that we have defined all algorithms for depth $P-1$. Fix $T_{0}\leq T_{1}$ with $T_{1}-T_{0}\geq 2^{P}$, and a labeled dataset $S$. We also fix a function $f_{S}\in\mathcal{F}$ realizing $S$ which will serve as base prediction function. First define $T_{1/2}:=\left\lfloor(T_{0}+T_{1})/2\right\rfloor$. This time divides the interval $(T_{0},T_{1}]$ in two epochs $(T_{0},T_{1/2}]$ and $(T_{1/2},T_{1}]$ of roughly equal length. Note that by construction, each epoch has length at least $2^{P-1}$. The algorithm proceeds separately on each epoch and more precisely aims to incorporate the information gathered on the first epoch $(T_{0},T_{1/2}]$ when making predictions during the second epoch $(T_{1/2},T_{1}]$. The procedure is detailed below.
@@
  • On the second epoch $(T_{1/2},T_{1}]$, the algorithm also performs a learning with expert advice algorithm but uses a larger set of experts, which will correspond to depth-$(P-1)$ R-Cover algorithms run with different datasets. Precisely, at the beginning of the epoch (before the prediction at time $t=T_{1/2}+1$), we consider all possible ways to label the queries $(x_{t})_{t\in(T_{0},T_{1/2}]}$ such that together with $S$ the increased dataset is still realizable by the function class $\mathcal{F}$. Formally, we enumerate all distinct datasets $S_{1},\ldots,S_{r}$ such that
    $\left\{S_{r^{\prime}},r^{\prime}\in[r]\right\}:=\left\{S^{\prime}=S\cup\left\{(x_{t},\tilde{y}_{t}),t\in(T_{0},T_{1/2}]\right\}:\tilde{y}_{t}\in\{0,1\},t\in(T_{0},T_{1/2}]\right\}\\ \cap\left\{S^{\prime}\subset\mathcal{X}\times\{0,1\}:\exists f\in\mathcal{F},\forall(x,y)\in S^{\prime},f(x)=y\right\}.$ (1)
[…]
For our purposes, we need a specific learning with expert advice algorithm A-Exp (see Algorithm 2). We defer its presentation to Section 3.2 for readability. This concludes the construction of the algorithm for depth $P$, horizon $T$, and dataset $S$, which is summarized in Algorithm 1.
[…]
@@ 3.2 Learning with expert advice algorithm
Instead of using the standard exponential weights algorithm [19] for learning with expert advice, we introduce a variant. We briefly recall the setup of prediction with $K$ experts and fixed horizon $T$ that is relevant to our present discussion. At each iteration $t\in[T]$, the environment chooses losses $\ell_{t,i}$ for each experts $i\in[K]$. The learner then selects an expert $\hat{i}_{t}\in[K]$ potentially randomly without knowledge of the losses at time $t$. Then, all losses at time $t$ are revealed to the learner and they incur the loss $\ell_{t,\hat{i}_{t}}$ of the selected expert. The goal of the learner is to minimize the regret:
$\text{Reg}(T):=\sum_{t=1}^{T}\ell_{t,\hat{i}_{t}}-\min_{i\in[K]}\sum_{t=1}^{T}\ell_{t,i}.$
The classical exponential weights forecaster or Hedge algorithm (see e.g. [19]) with parameter $\eta>0$ proceeds as follows. At time $t$, it computes the cumulative regret compared to each expert up to time $t$: $R_{t-1,i}:=\sum_{l=1}^{t-1}\ell_{l,\hat{i}_{l}}-\ell_{l,i}$ for all $i\in[K]$. It then randomly samples $\hat{i}_{t}\sim p_{t}$ where the distribution $p_{t}=(p_{t,i})_{i\in[K]}$ is defined via exponential weights $p_{t,i}:=e^{\eta R_{t-1,i}}/\sum_{j\in[K]}e^{\eta R_{t-1,j}}.$ We denote by $\mathcal{F}_{t}=(\ell_{l,i},l\leq t,i\in[K],\hat{i}_{l},l<t)$ the history up to time $t$ included. Hedge enjoys the following classical bound (see e.g. [19, Corollary 2.2]):
$\text{PReg}(T):=\sum_{t=1}^{T}\mathbb{E}_{\hat{i}_{t}}[\ell_{t,\hat{i}_{t}}\mid\mathcal{F}_{t}]-\min_{i\in[K]}\sum_{t=1}^{T}\ell_{t,i}\leq\frac{\ln K}{\eta}+\frac{T\eta}{2}.$
We will refer to the quantity on the left-hand side as the pseudo-regret $\text{PReg}(T)$. Using the standard choice of parameter $\eta=\sqrt{2\ln K/T}$, and assuming that the losses all have values in $[0,1]$, the previous equation directly gives an expected bound on the regret $\mathbb{E}[\text{Reg}(T)]\lesssim\sqrt{T\ln K}$. For our purposes, we need a refinement of this bound. Using [19, Theorem 2.1], we can obtain
$\text{PReg}(T)\leq\frac{\ln K}{\eta}+\frac{\eta}{2}\sum_{t=1}^{T}\sum_{i\in[K]}p_{t,i}r_{t,i}^{2},$ (2)
where $r_{t,i}:=\ell_{t,\hat{i}_{t}}-\ell_{t,i}$ is the instantaneous regret of the forecaster compared to expert $i$. Denote
@@
Eq. 2 yields a tighter bound than the standard regret bound if one selects $\eta\approx\sqrt{\ln K/\Delta_{T}}$ instead of the standard choice $\eta\approx\sqrt{\ln K/T}$. Achieving the corresponding bound without a prior knowledge of $\Delta_{T}$ can be easily performed via the standard doubling trick. Precisely, we use the exponential weights forecaster with initial parameter $\eta_{1}\approx\sqrt{2\ln K}$ until $\Delta_{t}\geq 1$, then restart the algorithm with a parameter $\eta_{2}\approx\eta_{1}/2$ until $\Delta_{t}\geq 4$. We continue the process by always restarting the algorithm with a quadrupled threshold for $\Delta$ and a corresponding parameter $\eta>0$ (roughly halved). The precise algorithm is given in Algorithm 2, which is the exponential weights forecaster variant that we use for our algorithm R-Cover. This variant enjoys the following regret bound, whose proof is given in Appendix B.
@@
3
4Initialization: $k=1$, $\Delta_{max,1}=1$, $\eta_{1}=\sqrt{2\ln K/(\Delta_{max,1}+1)}$, $R_{0,i}=0$ for all $i\in[K]$, $\Delta_{1}=0$
@@
6       Let $p_{t,i}=e^{\eta_{k}R_{t-1,i}}/\sum_{j\in[K]}e^{\eta_{k}R_{t-1,j}}$ and sample $\hat{i}_{t}\sim p_{t}$ independently from history
@@ 11            Reset $R_{t,i}=0$ for all $i\in[K]$, $\Delta_{k+1}=0$, and $k\leftarrow k+1$
Algorithm 2 Adaptive exponential weights forecaster A-Exp
@@
Suppose that all losses lie in $[0,1]$. Then, the pseudo-regret of the adaptive exponential weights forecaster A-Exp satisfies
$\textnormal{PReg}(T)\leq 8\sqrt{\max(\Delta_{T},1)\ln K},\quad T\geq 1.$
@@
$\textnormal{Reg}(T)\leq 12\sqrt{\max(\Delta_{T},1)\ln K}+2\ln\frac{1}{\delta}.$
@@ 4 Technical overview
As discussed above, the classification setting will be mostly sufficient to present our main proof ideas. Hence, in this section we mostly focus on this case.
@@ 4.1 A simple algorithm for a weaker regret guarantee
To motivate the form of R-Cover, we first consider a significantly simpler algorithm which essentially corresponds to R-Cover with depth $1$. In this simplest form, R-Cover subdivides the horizon $[T]$ into $K$ equal-length epochs and a learning with expert advice algorithm on each epoch on the projection of the function class $\mathcal{F}$ on query points $x_{t}$ from prior epochs. For instance, we can use the classical exponential weights forecaster algorithm (e.g. see [19]). This simplified algorithm which we call Cover is summarized in Algorithm 3.
@@
We formally prove this result in Appendix D. In this section, our goal is mostly to give key intuitions about the underlying strategy for the full algorithm R-Cover. To give some insights into why Cover already achieves sublinear regret, note that if the queries prior to some epoch $(T_{k-1},T_{k}]$ are “representative” of the queries during this epoch, then the cover $S_{k}$ constructed at the beginning of the epoch (line 3 of Algorithm 3) is a good representative set of relevant functions. Naturally, this holds if the underlying process $(x_{t})_{t\in[T]}$ is i.i.d.—that is $\sigma=0$. The crux of our analysis is to show that when the adversary is $\sigma$-smooth this still holds in an amortized sense: we show that the number of epochs for which prior queries $(x_{t})_{t\leq T_{k-1}}$ are not representative of the queries on the epoch $(T_{k-1},T_{k}]$ is bounded.
To quantify the notion of “representativeness”, we introduce the following quantity, which is essentially the maximum $\ell_{1}$ discrepancy between queries observed until some time $t_{0}<t$ and the query made at time $t$ on the function class $\mathcal{F}$. For any $0\leq t_{0}<t\leq T$, we define
@@
where $\mathcal{H}_{t-1}$ denotes all history available until the end of iteration $t-1$. One of our main contributions for the analysis of smoothed adversaries is the following result which bounds the number of epochs on which prior history is not representative.
@@
Let $T\geq 2$ and $\mathcal{F}:\mathcal{X}\to\{0,1\}$ be a function class with VC dimension $d$. Let $0=T_{0}<T_{1}<\ldots<T_{K}=T$ define epochs. Fix any parameters $q,\delta\in(0,1]$ and denote $w(T,\delta):=d\ln\left(\frac{T}{\sigma}\ln\frac{1}{\delta}\right)+\ln\frac{T}{\delta}+2$. Then, with probability at least $1-\delta$,
@@
for some universal constant $C\geq 1$. For a bound in expectation we can take $w(T):=d\ln\frac{T}{\sigma}+2$.
The proof of Proposition 7 uses some key results from [15]. At the high-level, we focus on a sub-sequence of $(x_{t})_{t\in[T]}$ by only keeping times which have a significant contribution $\gamma_{T_{k-1}}(t)\geq q$. Following arguments from [15], we next apply a bound (Lemma 21) to decouple the total sum of the terms $\gamma_{T_{k-1}}(t)$ via a tangent sequence to $(x_{t})_{t\in[T]}$ which is then bounded by the Wills functional of the function class $\mathcal{F}$ (Lemma 23). We defer a detailed sketch of proof to Appendix D.
@@
Proposition 7 shows that apart from $\tilde{\mathcal{O}}(1/(q\sigma))$ epochs, we pay at most a price $w(T,\delta)$ during each epoch $(T_{k-1},T_{k}]$ for the times $t\in(T_{k-1},T_{k}]$ when the cover constructed from queries prior to this epoch was not representative of query $x_{t}$ by some threshold $q$. Here, we view $w(T,\delta)$ as a reasonable price to pay on each epoch. Hence, intuitively, apart from $\tilde{\mathcal{O}}(1/(q\sigma))$ epochs, the cover constructed from queries on prior epochs is always representative up to threshold $q$.
@@
Up to the logarithmic factors, Proposition 7 is tight in the following sense. For any choice of the online mechanism to construct epochs and threshold $q$, a $\sigma$-smooth adversary can ensure that for $\mathcal{O}(1/(q\sigma))$ epochs $(T_{k-1},T_{k}]$, queries on prior epochs are not representative up to threshold $q$ from all times in $(T_{k-1},T_{k}]$. We detail below the scenarios for which Proposition 7 is tight.
@@
Because of the $\sigma$-smoothness constraint, the adversary cannot query the algorithm on completely different regions of the space $\mathcal{X}$ at each epoch. One possible strategy for the adversary is to switch distributions $\left\lfloor 1/\sigma\right\rfloor$ times during the learning process, possibly onto a completely new region of the space. This corresponds to $q=1$ in Proposition 7: at the start of $\left\lfloor 1/\sigma\right\rfloor$ epochs $(T_{k-1},T_{k}]$, the adversary switches query distributions $\mu_{t}$ and selects a distribution with support on a new region for which prior queries are irrelevant. This results in $\gamma_{T_{k-1}}(t)=1$ for all $t\in(T_{k-1},T_{k}]$.
@@
A more refined strategy for the adversary is to select a parameter $q$ and at the start of a new epoch $(T_{k-1},T_{k}]$, switch the query distribution as follows. They construct a new mixture distribution $\mu_{k}:=q\nu_{k}+(1-q)\mu_{0}$ where with probability $q$ the learner is queried on a new distribution $\nu_{k}$ with say completely new support compared to the history, and with probability $1-q$ the learner is queried on a base measure $\mu_{0}$ that is very similar to previous queries. This results in $\gamma_{T_{k-1}}(t)\geq q$ for all $t\in(T_{k-1},T_{k}]$. On one hand, during the epoch, the adversary could only test the learner on a fraction $q$ of “truly adversarial” queries sampled from $\mu_{k}$. On the other hand, the smoothness constraint is now easier to satisfy and we can check that the adversary can afford to corrupt $\approx 1/(q\sigma)$ epochs. This precisely corresponds to the bound from Proposition 7 up to logarithmic factors. As it turns out, this mixture strategy is stronger for the adversary and choosing $q\approx 1/\sqrt{T}$ is the strategy that yields the lower bound from Theorem 3.
@@
The statement from Proposition 7 is written specifically for classification, for which analyzing the $\ell_{1}$ diameter as defined in $\gamma_{t_{0}}(t)$ in Eq. 3 is amenable. The proof of Proposition 7 requires controlling the complexity of the class $\{\mathbbm{1}[f\neq g]:f,g\in\mathcal{F}\}$ which has VC dimension bounded by $2d$ if $\mathcal{F}$ has VC dimension $d$. While the VC dimension behaves nicely with this self-difference operation, this is not the case for the fat-shattering dimension which is known to behave somewhat wildly with the addition [1].222[1] notes that the function class $\mathcal{F}$ of increasing functions on $[0,1]$ always has fat-shattering dimension one at any scale, while =$\mathcal{F}-\mathcal{F}=\{f-g:f,g\in\mathcal{F}\}$ has infinite shattering dimension at all scales. For the regression setting, we need to localize this difference class around an oblivious benchmark function $f^{\star}$. The localized analog of $\gamma_{t_{0}}(t)$ that we use in our proofs is defined in Eq. 17. The corresponding generalization of Proposition 7 is Lemma 11. In this general regression setting, the term in $w(T,\delta)$ from Proposition 7 depending on the VC dimension $d$ is replaced by the Wills functional of $\mathcal{F}$.
@@
for some constant $C\geq 1$, where we used Jensen’s inequality in the last inequality. Next, for each epoch $k\in[K]$, denote by $f_{k}\in S_{k}$ the function in the cover that had the correct labeling for $f^{\star}$, that is, $f_{k}(x_{t})=f^{\star}(x_{t})$ for all $t\in[T_{k-1}]$. Because $f_{k}$ is one of the experts considered during epoch $k$, it suffices to bound the remaining term